Let g be a function from Z+ (the set of positive integers) to Q (the set of rational numbers) defined by (x, y) element of g iff y = (g is a subset of Z+ mapped with Q) and let f be a function on Z + defined by (x, y) element of f iff y = 5x2 + 2x – 3 (f subset of Z+ mapped with Z+)

Which one of the following statements regarding the function g is TRUE? 

(Remember, g is a subset of Z+ mapped with Q.)

1. g can be presented as a straight line graph.

2. g is injective.

3. g is surjective.

4. g is bijective.



1
Expert's answer
2021-07-20T12:36:20-0400

Let "g" be a function from "Z^+" (the set of positive integers) to "Q" (the set of rational numbers) defined by


"(x, y)\in g\ \text{iff } y=4x-\dfrac{3}{7}\ (g\sube Z^+\times Q)"

and let "f" be a function on "Z^+" defined by


"(x, y)\in g\ \text{iff } y=5x^2+2x-3\ (f\sube Z^+\times Z^+)"



We consider the statements provided in the different alternatives:

1. "g" is not defined on the set of real numbers thus "g" cannot be depicted as a straight line graph.

Only positive integers can be present in the domain of "g."

It is the case that ordered pairs such as "(1, 3\dfrac{4}{7}), (2, 7\dfrac{4}{7}), (3,11\dfrac{4}{7}),..." belong to "g" and these pairs can be presented as dots in a graph. 


2. We prove that "g" is indeed injective:


"g(u)=g(v)=>4u-\dfrac{3}{7}=4v-\dfrac{3}{7}"

"=>4(u-v)=0=>u=v"

"g" is injective.


3. The function "g" is NOT surjective. 

Counterexample


"y=\dfrac{1}{2}, \dfrac{1}{2}\in Q"

Then


"y=4x-\dfrac{3}{7}=\dfrac{1}{2}=>4x=\dfrac{13}{14}=>x=\dfrac{13}{56}"

"\dfrac{13}{56}\not\in Z^+"

We conclude that "g" is not surjective. 


4. Since "g" is not surjective, then "g" is not bijective.  



Answer

2. "g" is injective.